OpenStudy (sushi121212):

The areas of 2 circles are 36π snd 64π. What is a ratio of the diameters? Of the circumcenters?

jhonyy9 (jhonyy9):

ratio of diameters 36pi mean r=6 so diameter 2*6=12 64pi mean r=8 so diameter 2*8=16 ratio 12/16 = 3/4

OpenStudy (jdoe0001):

hmmm can you follow what jhonyy9 suggested ?

OpenStudy (sushi121212):

Does it mean that the r is the square root of the numbers 36 and 64? What about the circumferences?